# Prior-Guided Diffusion Planning for Offline Reinforcement Learning

### v1

@@ C.3 Statistical Details
$\bar{\sigma}_{\text{avg}}=\sqrt{\sum_{i=1}^{n}\bar{\sigma}_{i}^{2}}\bigg{/}n,\quad\text{where}\leavevmode\nobreak\ \bar{\sigma}_{i}=\frac{\sigma_{i}}{\sqrt{k}},$
@@ Appendix D Diffusion Veteran (DV)
Input: Dataset $D$, Planning horizon $H$, Planning stride $m$, Candidate num $N$
@@
1:Compute discounted returns: $R^{(\tau)}=\sum_{h=0}^{\infty}\gamma^{h}r^{(\tau+h)}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{for every step }\tau$
@@
12:     Randomly generate $N$ candidate trajectories using $g_{\mathbf{s}}$ while fixing the first state as $\mathbf{s}^{(\tau)}$
13:     Select the trajectory with the highest estimated value $V$ among the $N$ candidates
@@
Table 3: Normalized scores on the D4RL benchmark. we compare Prior Guidance (PG) with the original implementation DV and DV* under different $N$s. We compute the mean and the standard error over 5 random seeds.
@@ Appendix E Full results of Table 1
BC CQL IQL SfBC DQL IDQL QGPO SRPO Diffuser AD DD HD DV* PG
Walker2d-M 6.6 79.2 78.3 77.9$\pm$2.5 87.0$\pm$0.9 82.5 86.0$\pm$0.7 84.4$\pm$1.8 79.6$\pm$0.6 84.4$\pm$2.6 82.5$\pm$1.4 84.0$\pm$0.6 79.5$\pm$0.3 82.3$\pm$0.2
Walker2d-M-R 11.8 26.7 73.9 65.1$\pm$5.6 95.5$\pm$1.5 85.1 84.4$\pm$4.1 84.6$\pm$2.9 70.6$\pm$1.6 84.7$\pm$3.1 75.0$\pm$4.3 84.1$\pm$2.2 83.5$\pm$2.5 83.7$\pm$1.0
Walker2d-M-E 6.4 111.0 109.6 109.8$\pm$0.2 110.1$\pm$0.3 112.7 110.7$\pm$0.6 114.0$\pm$0.9 106.9$\pm$0.2 108.2$\pm$0.8 108.8$\pm$1.7 107.1$\pm$1.1 109.0$\pm$0.1 109.4$\pm$0.1
Hopper-M 29.0 58.0 66.3 57.1$\pm$4.1 90.5$\pm$4.6 65.4 98.0$\pm$2.6 95.5$\pm$0.8 74.3$\pm$1.4 96.6$\pm$2.7 79.3$\pm$3.6 99.3$\pm$0.3 84.1$\pm$2.5 97.5$\pm$0.6
Hopper-M-R 11.3 48.6 94.7 86.2$\pm$9.1 101.3$\pm$0.6 92.1 96.9$\pm$2.6 101.2$\pm$0.4 93.6$\pm$0.4 92.2$\pm$1.5 100.0$\pm$0.7 94.7$\pm$0.7 91.3$\pm$0.1 91.3$\pm$0.3
Hopper-M-E 111.9 98.7 91.5 108.6$\pm$2.1 111.1$\pm$1.3 108.6 108.0$\pm$2.5 100.1$\pm$5.7 103.3$\pm$1.3 111.6$\pm$2.0 111.8$\pm$1.8 115.3$\pm$1.1 109.9$\pm$1.0 110.4$\pm$0.1
HalfCheetah-M 36.1 44.4 47.4 45.9$\pm$2.2 51.1$\pm$0.5 51.0 54.1$\pm$0.4 60.4$\pm$0.3 42.8$\pm$0.3 44.2$\pm$0.6 49.1$\pm$1.0 46.7$\pm$0.2 50.9$\pm$0.1 45.6$\pm$0.5
HalfCheetah-M-R 38.4 46.2 44.2 37.1$\pm$1.7 47.8$\pm$0.3 45.9 47.6$\pm$1.4 51.4$\pm$1.4 37.7$\pm$0.5 38.3$\pm$0.9 39.3$\pm$4.1 38.1$\pm$0.7 46.4$\pm$0.2 46.4$\pm$0.4
HalfCheetah-M-E 35.8 62.4 86.7 92.6$\pm$0.5 96.8$\pm$0.3 95.9 93.5$\pm$0.3 92.2$\pm$1.2 88.9$\pm$0.3 89.6$\pm$0.8 90.6$\pm$1.3 92.5$\pm$0.3 92.3$\pm$0.6 95.2$\pm$0.1
@@ Appendix G Extensive Results
walker2d-medium-v2 81.3$\pm$0.2 82.3$\pm$0.2
walker2d-medium-replay-v2 83.2$\pm$2.1 83.7$\pm$1.0
walker2d-medium-expert-v2 109.6$\pm$0.2 109.4$\pm$0.1
hopper-medium-v2 96.9$\pm$1.4 97.5$\pm$0.6
hopper-medium-replay-v2 94.5$\pm$0.7 91.3$\pm$0.3
hopper-medium-expert-v2 110.4$\pm$0.1 110.4$\pm$0.1
halfcheetah-medium-v2 54.7$\pm$0.2 45.6$\pm$ 0.5
halfcheetah-medium-replay-v2 46.6$\pm$0.1 46.4$\pm$0.4
halfcheetah-medium-expert-v2 95.1$\pm$0.2 95.2$\pm$0.1
@@
antmaze-medium-play-v2 87.0$\pm$1.5 87.8$\pm$1.5
antmaze-medium-diverse-v2 89.6$\pm$1.8 87.3$\pm$1.7
antmaze-large-play-v2 82.2$\pm$2.4 82.4$\pm$2.1
antmaze-large-diverse-v2 77.0$\pm$3.2 76.0$\pm$1.9
@@
maze2d-umaze-v1 137.1$\pm$1.5 139.2$\pm$1.3
maze2d-medium-v1 159.4$\pm$3.5 159.5$\pm$0.8
maze2d-large-v1 196.2$\pm$2.6 195.2$\pm$2.3
@@
walker2d-medium-v2 81.3$\pm$0.2
walker2d-medium-replay-v2 82.5$\pm$0.2
walker2d-medium-expert-v2 108.7$\pm$0.1
hopper-medium-v2 88.9$\pm$2.3
hopper-medium-replay-v2 96.2$\pm$0.0
hopper-medium-expert-v2 110.5$\pm$0.1
halfcheetah-medium-v2 44.3$\pm$2.1
halfcheetah-medium-replay-v2 43.0$\pm$0.2
halfcheetah-medium-expert-v2 84.9$\pm$0.5
@@
antmaze-medium-play-v2 87.2$\pm$1.2
antmaze-medium-diverse-v2 93.0$\pm$1.4
antmaze-large-play-v2 83.4$\pm$1.4
antmaze-large-diverse-v2 79.4$\pm$1.0
@@ Appendix H Ablation Stuides
We conducted ablation studies on the behavior regularization coefficient $\alpha$ of PG across MuJoCo, Kitchen, AntMaze, and Maze2D tasks. The normalized scores of PG for different values of $\alpha$ are reported in Table 111213 and 14.
@@
walker2d-medium-v2 77.2$\pm$0.6 81.4$\pm$0.3 82.3$\pm$0.2 74.5$\pm$0.8 70.8$\pm$2.2
walker2d-medium-replay-v2 61.9$\pm$1.0 83.7$\pm$1.2 78.3$\pm$0.8 52.3$\pm$1.5 50.6$\pm$1.4
walker2d-medium-expert-v2 109.4$\pm$0.1 109.1$\pm$0.1 108.6$\pm$0.2 108.4$\pm$0.1 108.5$\pm$0.2
hopper-medium-v2 93.5$\pm$0.4 97.5$\pm$0.6 68.9$\pm$0.8 54.3$\pm$0.5 51.4$\pm$0.9
hopper-medium-replay-v2 91.3$\pm$0.3 90.8$\pm$0.7 89.8$\pm$0.2 61.4$\pm$0.3 34.8$\pm$0.8
hopper-medium-expert-v2 110.4$\pm$0.1 110.3$\pm$0.1 110.3$\pm$0.3 87.0$\pm$0.9 68.2$\pm$1.3
halfcheetah-medium-v2 45.6$\pm$0.5 40.2$\pm$0.6 43.4$\pm$0.3 43.4$\pm$0.3 43.1$\pm$0.4
halfcheetah-medium-replay-v2 46.4$\pm$0.4 45.0$\pm$0.2 44.1$\pm$0.4 43.0$\pm$0.5 39.0$\pm$0.5
halfcheetah-medium-expert-v2 8.2$\pm$1.1 95.2$\pm$0.1 94.9$\pm$0.2 95.1$\pm$0.2 92.8$\pm$0.4
@@
antmaze-medium-play-v2 47.0$\pm$4.2 87.8$\pm$1.5 83.8$\pm$1.6 84.0$\pm$1.3 80.2$\pm$1.5
antmaze-medium-diverse-v2 86.0$\pm$1.4 87.3$\pm$1.7 80.3$\pm$1.8 72.1$\pm$2.5 68.4$\pm$2.8
antmaze-large-play-v2 34.5$\pm$4.8 74.6$\pm$2.5 76.8$\pm$2.0 71.3$\pm$1.7 82.4$\pm$2.1
antmaze-large-diverse-v2 21.8$\pm$3.2 70.4$\pm$1.8 69.7$\pm$1.8 76.0$\pm$1.9 68.5$\pm$2.1
@@
maze2d-umaze-v1 113.3$\pm$2.5 91.9$\pm$3.7 126.3$\pm$1.7 139.2$\pm$1.3 122.2$\pm$2.1
maze2d-medium-v1 155.6$\pm$0.8 121.1$\pm$4.8 159.5$\pm$0.8 148.4$\pm$1.7 152.1$\pm$1.2
maze2d-large-v1 143.4$\pm$5.1 151.3$\pm$4.7 195.2$\pm$2.3 185.0$\pm$5.3 183.9$\pm$4.8
@@
walker2d-medium-v2 82.3$\pm$0.2 82.0$\pm$0.3 78.6$\pm$1.4
walker2d-medium-replay-v2 83.7$\pm$1.0 83.7$\pm$0.4 82.4$\pm$0.3
walker2d-medium-expert-v2 109.4$\pm$0.1 109.1$\pm$0.0 103.5$\pm$0.2
hopper-medium-v2 97.5$\pm$0.6 96.5$\pm$0.9 97.0$\pm$0.1
hopper-medium-replay-v2 91.3$\pm$0.3 93.5$\pm$0.0 92.4$\pm$0.0
hopper-medium-expert-v2 110.4$\pm$0.1 110.3$\pm$0.0 111.5$\pm$0.0
halfcheetah-medium-v2 45.6$\pm$ 0.5 48.8$\pm$0.5 43.5$\pm$0.1
halfcheetah-medium-replay-v2 46.4$\pm$0.4 46.0$\pm$0.1 44.9$\pm$0.1
halfcheetah-medium-expert-v2 95.2$\pm$0.1 95.4$\pm$0.1 92.5$\pm$0.1
@@
antmaze-medium-play-v2 87.8$\pm$1.5 87.6$\pm$2.2 76.3$\pm$2.5
antmaze-medium-diverse-v2 87.3$\pm$1.7 86.3$\pm$2.0 80.0$\pm$2.3
antmaze-large-play-v2 82.4$\pm$2.1 79.7$\pm$1.8 85.2$\pm$1.5
antmaze-large-diverse-v2 76.0 $\pm$ 1.9 77.0$\pm$1.9 70.9$\pm$2.2
@@
maze2d-umaze-v1 139.2$\pm$1.3 136.2$\pm$2.1 132.0$\pm$2.2
maze2d-medium-v1 159.5$\pm$0.8 150.4$\pm$1.6 170.8$\pm$1.0
maze2d-large-v1 195.2$\pm$2.3 186.6$\pm$2.3 193.2$\pm$2.0

### v2

@@ C.3 Statistical Details
$\bar{\sigma}_{\text{avg}}=\sqrt{\sum_{i=1}^{n}\bar{\sigma}_{i}^{2}}\bigg/n,\quad\text{where}\penalty 10000\ \bar{\sigma}_{i}=\frac{\sigma_{i}}{\sqrt{k}},$
@@ Appendix D Diffusion Veteran (DV)
Input: Dataset $D$, Planning horizon $H$, Planning stride $m$, Candidate num NN
@@
1:Compute discounted returns: $R^{(\tau)}=\sum_{h=0}^{\infty}\gamma^{h}r^{(\tau+h)}\penalty 10000\ \penalty 10000\ \text{for every step }\tau$
@@
12:  Randomly generate NN candidate trajectories using $g_{\mathbf{s}}$ while fixing the first state as $\mathbf{s}^{(\tau)}$
13:  Select the trajectory with the highest estimated value $V$ among the NN candidates
@@
Table 3: Normalized scores on the D4RL benchmark. we compare Prior Guidance (PG) with the original implementation DV and DV* under different NNs. We compute the mean and the standard error over 5 random seeds.
@@ Appendix E Full results of Table 1
BC CQL IQL SfBC DQL IDQL QGPO SRPO Diffuser AD $D$ HD DV* PG
[…]
@@ Appendix G Extensive Results
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ Appendix H Ablation Stuides
We conducted ablation studies on the behavior regularization coefficient $\alpha$ of PG across MuJoCo, Kitchen, AntMaze, and Maze2D tasks. The normalized scores of PG for different values of $\alpha$ are reported in Table 11, 12, 13 and 14.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Appendix I Why PG underperform on standard MuJoCo tasks?
Table 16: Performance comparison across standard MuJoCo tasks.
DQL DV* PG (planner) PG (policy)
walker2d-medium-v2 87.0 79.5 82.3 84.5 $\pm$ 0.26
walker2d-medium-replay-v2 95.5 83.5 83.7 96.4 $\pm$ 2.49
walker2d-medium-expert-v2 110.1 109.0 109.4 109.8 $\pm$ 0.23
hopper-medium-v2 90.5 84.1 97.5 98.4 $\pm$ 1.01
hopper-medium-replay-v2 101.3 91.3 91.3 100.5 $\pm$ 1.27
hopper-medium-expert-v2 111.1 109.9 110.4 111.4 $\pm$ 0.57
halfcheetah-medium-v2 51.1 50.9 45.6 54.7 $\pm$ 0.42
halfcheetah-medium-replay-v2 47.8 46.4 46.4 48.6 $\pm$ 0.34
halfcheetah-medium-expert-v2 96.8 92.3 95.2 95.5 $\pm$ 0.26
Average 87.9 83.0 84.6 88.9
PG is a diffusion planner-based method designed for long-horizon planning, where capturing temporally extended dependencies is essential. However, MuJoCo locomotion tasks involve short-horizon, dense-reward control focused on making agents run faster, which does not require such lookahead capabilities [19]. As a result, diffusion planner-based methods, including PG, often underperform compared to diffusion policy-based approaches in these settings, as shown in Table 1. While the additional structure and trajectory-level modeling of planners benefit challenging long-horizon domains such as AntMaze, Kitchen, or Maze2D, they provide little advantage and can even be slightly detrimental in simpler environments where policies can already achieve near-optimal performance.
To examine whether the observed performance gap in MuJoCo tasks from the use of the diffusion planner rather than a limitation of the PG framework itself, we conduct an additional experiment by applying PG to a diffusion policy. Specifically, we replaced the standard Gaussian prior of a behavior-cloned diffusion policy with a learnable, behavior-regularized prior, following the same training procedure used in PG.
Table LABEL:tab:pg_policy shows that replacing the diffusion planner in PG with a diffusion policy (PG policy) consistently improved performance across MuJoCo tasks, raising the average score from 84.6 (PG planner) to 88.9 and surpassing the DQL baseline score of 87.9. This demonstrates that PG remains effective even in short-horizon environments, suggesting that the lower performance of PG in MuJoCo tasks is due to the nature of trajectory-level planning rather than a limitation of the method itself. The results for PG (policy) are averaged over 3 random seeds and reported as the mean $\pm$ standard error.